Find the vertex of f(x)=x^2+5x+4 step by step.
Answer by Alan3354(69443)

Find the vertex of f(x)=x^2+5x+4 step by step.
----------
The vertex is on the LOS, Line of Symmetry, x = -b/2a
---
x = -5/2 is the LOS
f(-5/2) = 25/4 - 25/2 + 4 = -9/4
--> vertex @ (-5/2,-9/4)